OpenStudy (anonymous):

can someone help me with this 1 + sec^2(x) sin^2(x) = sec^2(x)

OpenStudy (anonymous):

prove the equation

OpenStudy (haseeb96):

1+ sec^2 (x) [1 +cos^2 (x)] = sec^2 (x) 1+ sec^2(x) + sec^2 (x) .cos^2 (x) = sec^2 (x) 1+ sec^2(x) =sec^2 (x) is yur question correct?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

sin^2(x) doesnt equal 1+cos^2 (x) it equals 1-cos^2(x)

OpenStudy (unklerhaukus):

1 + sec^2 . sin^2 = sec^2 1 + 1/cos^2 . sin^2 = sec^2 1 + sin^2/cos^2 = sec^2 1 + tan^2 = sec^2

OpenStudy (anonymous):

thanks dude

OpenStudy (unklerhaukus):

Which is an identity. _ _ _ [if you want to do it another way] 1 + sec^2(x) sin^2(x) = sec^2(x) 1 + 1/cos^2 . sin^2 =1/cos^2 cos^2 + sin^2 = 1 Which might be a more familiar identity .

OpenStudy (anonymous):

convert sec^2(x) into 1/cos^2(x) 1+tan^2(x) Hence proved 1+tan^2(x) = sec^2(x) Any questions ?

OpenStudy (unklerhaukus):

@Haseeb96 's method would have worked too, 1+ sec^2 [1 - cos^2] = sec^2 1+ sec^2 - cos^2sec^2 = sec^2 1+ sec^2 - cos^2/cos^2 = sec^2 1+ sec^2 - 1 = sec^2 sec^2 = sec^2

OpenStudy (anonymous):

thanks